OpenStudy (anonymous):

use system of linear equations: at a local high school basketball game, 1435 tickets were sold. a student admission ticket cost 1.50 and an adult admission ticket cost 5.00. The total ticket receipts for the game were 3552.50. how many of each type of ticket were sold?

OpenStudy (anonymous):

x=numbers of student tickets y=number of adult tickets total cost of student admission = 1.5x and total cost of adult admission=5y leaves you with the two equations x+y=1435 1.5x+5y=3552.50 rearranging your first equation y=1435-x then you substitute that into your 2nd equation 1.5x+5(1435-x)=3552.5 and solve for x 1.5x-5x+7175=3552.5 -3.5x=-3622.5 x=1035 = number of student tickets sold 1435-1035=y=400=number of adult tickets sold